OpenStudy (anonymous):

if r and s are solutions of x^2 + 6x - 2 = 0, what is the value of r^3 + s^3.

OpenStudy (anonymous):

guess there is no choice to to actually solve

OpenStudy (anonymous):

Yeah try using the quadratic formula to find r and s

OpenStudy (campbell_st):

funny how the same questions keep popping up

OpenStudy (blockcolder):

If r and s are the solutions, then r+s=-6 and rs=-2. \[r^3+s^3=(r+s)(r^2-rs+s^2)=(r+s)((r+s)^2-3rs)=(-6)(36+6)=-252\]

OpenStudy (anonymous):

cool, a snap way! but i can't see it all

OpenStudy (anonymous):

What does the addition of the two zeroth solutions represent?

OpenStudy (anonymous):

Is there something special about it?

OpenStudy (blockcolder):

In a quadratic equation ax^2+bx+c=0, the solutions add up to -b/a while their product is c/a.

OpenStudy (anonymous):

sum of roots is \(-\frac{b}{a}\), product of roots is \(\frac{c}{a}\) and that cool trick was to write the sum of the cubes in terms of those. i will have to try to remember this

OpenStudy (anonymous):

I never learned this. Wow

OpenStudy (anonymous):

Cool trick

OpenStudy (anonymous):

i never would have thought of \[r^3+s^3=(r+s)(r^2-rs+s^2)=(r+s)((r+s)^2-3rs)\]

OpenStudy (blockcolder):

Ah, the wonders of algebra. =))

OpenStudy (anonymous):

thx you were a big help

OpenStudy (blockcolder):

No problem. :D